OpenStudy (anonymous):

which of the following statements are true? A. P(A and B) = P(A) * P(B|A) B. P(A and B) = P(B) * P(A|B) C. P(A|B)= P(B) if A and B are independent D. P(A|B)= P(A) if A and B are independent

OpenStudy (anonymous):

first one is true for sure

OpenStudy (anonymous):

as is the second one switch A and B and they are identical

OpenStudy (anonymous):

P(A|B)= P(A) is the definition of independence, so that is true as well

OpenStudy (anonymous):

only C is false, in fact it is rather silly

OpenStudy (anonymous):

Okay thank you! Could you briefly explain why C is false and D is correct?

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

first of all \[P(A|B)\] which reads "the probability of A given B" is the probability that A occurs if you know B has occurred if knowing B has occurred gives no information about A occurring, i.e. if \[P(A|B)=P(A)\] the you say A and B are "independent" that is the definition of independence

OpenStudy (anonymous):

now this is not at all the same as \[P(A|B)=P(B)\] you can see first of all that they look different

OpenStudy (anonymous):

Oh I see, thank you! I was confused about that for a while thanks

OpenStudy (anonymous):

yw i can give an example if you like

OpenStudy (anonymous):

sure all the more to help me

OpenStudy (anonymous):

are you familiar with dice? that is always a good example

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

give me a second

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

lets do this a simpler way suppose \[P(A)=\frac{2}{3}\], A and B are independent so \[P(A|B)=\frac{2}{3}\] as well by definition \[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

but this does not mean that \(P(B)\) must also be \(\frac{2}{3}\) it just means the ratio must be \(\frac{2}{3}\) so for example it could be \[\frac{P(A\cap B)}{P(B)}=\frac{\frac{2}{7}}{\frac{3}{7}}\] making \[P(B)=\frac{3}{7}\]

OpenStudy (anonymous):

we could do an example with a venn diagram as well

OpenStudy (anonymous):

I think I understand, Thank you so much!

OpenStudy (anonymous):

yw try this put \(P(A)=.8, P(B)=.5, P(A\cap B)=.4\) show that \(A\) and \(B\) are independent